# Hybrid Many-Objective Optimization in Probabilistic Mission Design for Compliant and Effective UAV Routing

### v1

@@ 3.2. Statistical Relational Maps for Reasoning on Airspace Regulations
We have recently introduced Statistical Relational Maps (StaR Maps) to encapsulate uncertain environments made up of semantically annotated features, i.e., geometry that has been assigned error parameters and labeled with descriptive tags (Flade et al., 2024). Rather than containing a graphical representation of the environment, a StaR Map parameterizes hybrid probabilistic relations between points in state-space and sets of features.
@@
According to StaR Maps, a spatial relation with domain and codomain of describes a categorical relationship between a point in , a tag of map features in and a map in . For example, this may categorize points into whether they lie within or over a feature with the respective tag. Analogously, a relation with domain and codomain of describes a quantitative relationship, e.g., the distance between a point and the closest element in the set of features with the respective tag. The set itself depends on the application, e.g., in an autonomous driving context might contain references to roads, crossings, or traffic lights.
@@
As the evaluation of a spatial relation will differ for each perturbed map, StaR Maps computes the statistical moments to fit the distributions chosen for each relation. To estimate the parameters for a specific spatial relation and a type on a point , we compute . With the set at hand, we obtain the statistical moments, e.g., mean and variance, of . Using maximum likelihood estimation or moment matching with the desired distribution, for example, a Bernoulli distribution for - or a Gaussian distribution for -domain relations, we store the resulting parameters within a StaR Map. The so-obtained spatial relations can then be compiled into distributional ground atoms in probabilistic first-order logic, which is how we proceed in Section 3.4.
@@ 3.3. Physical Objectives
In practice, one may require many objectives for UAVs to pursue. Here, we consider exemplary physical and geo-referenced objective function formulations evaluating a path regarding the risk of injury for city residents, noise emission, radio signal disturbance, and energy consumption, which have been presented in detail in a prior publication (Hohmann et al., 2023). Hence, in the following, we re-introduce these objective functions briefly and refer to the original work for an in-depth discussion. We categorize and discuss these objectives in two forms: grid-based and non-grid-based.
@@ 3.4. The Constitutional Language
Programming an agent’s Constitution: Here, we show an example Constitution where both a simple model for local airspace regulations and a set of physical objectives is imposed on the UAV’s path planning. Note that the choice of parameters is not done within the Constitution but at runtime through the operator or automatically selected by maximizing a mission’s probability of satisfying all requirements. {minted} [ frame=none, autogobble, fontsize=, xleftmargin=20pt, linenos ]python # Mission setting and parameter options star_map(”./environments/example.star”). parameter regular_license, special_license. parameter day, night.
# Further knowledge and rules take_off_mass normal(20.0, 1.0). light_drone if take_off_mass ¡ 5.0. field line_of_sight if day and distance(pilot) ¡ 100. field airspace if line_of_sight and regular_license and take_off_mass ¡ 10.0 or special_license and take_off_mass ¡ 25.0.
# The UAV’s objectives field objective airspace. field objective radio(”./models/radio.py”). path objective energy(”./models/energy.py”).
@@
A Constitution written in CoLa as exemplified in Listing 3.4 is written in three parts: First, one defines the setting by referring to a StaR Map, providing the probabilistic spatial relations one may use to encode the legal requirements, and determining the parameters under which the mission may be carried out, e.g., weather conditions, licensing information, or abstract UAV properties. Second, one defines further background knowledge and rules on which the compliance of an agent’s state can be reasoned. Third, one defines which quantities shall be treated as objectives across the navigation space or on a per-path basis.
While on the one hand, physical objectives as discussed in Section 3.3 may be incorporated through their mathematical description in a language such as Python 333https://www.python.org/, the logical reasoning on airspace regulations is facilitated by compiling the relevant parts of the Constitution into a probabilistic logic programming language supporting hybrid probabilistic inference, i.e., on discrete and continuous distributions simultaneously, such as Hybrid ProbLog (Nitti et al., 2016). Other systems with more advanced neuro-symbolic capabilities such as DeepProbLog (Manhaeve et al., 2018), NeurASP (Yang et al., 2020) or SLASH (Skryagin et al., 2023) are left for future work, e.g., to integrate end-to-end learning with deep learning models.
@@
To query for the probability that the Constitution’s modeled requirements are satisfied for a location , it needs to be compiled into the chosen backend, such as Hybrid ProbLog. This is achieved by (i) replacing syntactic elements, e.g., conjunctions (and) are written as comma-separated lists in ProbLog, and (ii) inserting the parameters of the StaR Map. We refer to the original paper for details on the latter part (Flade et al., 2024).
Such a probabilistic logic program must first be solved, i.e., the models under which it is satisfied must be found to perform probabilistic inference. While details will differ across off-the-shelf solving pipelines, the rough process can be described in two steps. First, the program is grounded, meaning all variables are substituted, and the program is rewritten in Conjunctive Normal Form (CNF). Second, the actual solver consumes the CNF, producing a Disjunctive Normal Form (DNF) where every conjunction represents a possible assignment of ground atoms under which the (constrained for the currently queried objective) program is satisfied. For instance, one may employ clingo (Gebser et al., 2017), a state-of-the-art suit for solving programmatic first-order logic based on Answer Set Programming (Brewka et al., 2011).
For exact probabilistic inference, one needs to assign the probabilities of atom to take on the value assigned by model . These parameters have been either written down directly in the Constitution, e.g., the take-off-mass in Listing 3.4, or are provided by the StaR Map. In turn, the probability of the Constitution being satisfied given the state and measurement at time is then obtained via the sum-product
@@ 3.5. Many-Objective Path Planning
We facilitate the search for a Pareto-set of paths considering the objectives stated in a CoLa program through, e.g., the physical objectives defined in Section 3.3 and airspace-dependent regulations stated in probabilistic first-order logic by employing the Versatile Intelligent Aerial Streets (VIAS) framework for many-objective path planning as presented in a prior publication (Hohmann et al., 2021). Besides the reference implementation of the overall architecture (see Figure 1) being made available as an open-source package, VIAS is also available as a stand-alone solution.
In its first stage, VIAS applies the single-objective Dijkstra algorithm to weighted aggregations of all grid-based objective functions. This results in three-dimensional polygonal paths guaranteed optimal in the employed grid structure. Next, VIAS smooths and approximates the polygonal paths and obtains three-dimensional NURBS curves (Piegl and Tiller, 1997a, b). These splines are henceforth treated as initial solutions in the second stage, a many-objective evolutionary optimization problem calculating an -dimensional Pareto-set of smooth paths that can be thought of as a set of optimal trade-off solutions between all considered objective functions.
@@
The smoothing and approximation between the two optimization stages is crucial. At this point, the path representation changes from a polygonal path definition to a NURBS definition. Here we will only provide a brief discussion; the interested reader is referred to (Hohmann et al., 2023). VIAS first smoothes the polygonal path by applying a Gaussian filter with kernel . Then, the smoothed paths are approximated in the form of a NURBS curve. For this operation, the critical parameter is the number of control points, , for the resulting NURBS curve. While a low number may lead to an excessive approximation error, a large number increases the size of the optimization vector, expanding the search space during the evolutionary optimization step.
@@ 3.6. Mission Clearance
Given a CoLa program and the subsequently obtained Pareto-set of paths from VIAS as outlined in Sections 3.4 and 3.5 respectively, one may still wonder if the results are overall sufficiently likely to satisfy the modeled airspace regulations. Hence, based on our prior publication on Clearance, Explanation, and Optimization in ProMis (Kohaut et al., 2024), we discuss each of these steps with regard to Pareto-optimal proposal paths.
@@
Note that this Clearance does not necessarily coincide with a path being the knee-point or even optimal with regard to the airspace requirements. On the other hand, an extreme-point of the Pareto-set for an unrelated objective may still be granted Clearance, i.e., this step has to be performed on top of the many-objective routing itself.
@@ 3.7. Mission Explanation
Consider again the example CoLa program in Listing 3.4. Here, two sets of parameters were defined, one on the employed license and one on the time of day. With regard to granting clearance, it is easy to see in this case that a special operations license allows for easier Clearance since no requirement on visual line-of-sight is imposed.
@@ 3.8. Mission Optimization
| ProMis | Legal Model | Listing 4.1 | |
@@ 4. Experiments
Throughout Section 3, we have presented an approach for effective and compliant UAV routing under many objectives. In this Section, we evaluate our methods on real-world, crowd-sourced data from the city of Paris, demonstrating how our framework not only jointly models airspace regulations and physical objectives but also produces suitable Pareto-sets trading-off the objectives and applying Clearance, Explanation, and Optimization on the mission setting to further improve on the found solutions.
@@ 4.1. Scenario Definition
For our experiments, we consider a large-scale urban environment with data sourced from OpenStreetMap (OpenStreetMap contributors, 2017) and OpenCellID (OpenCellid contributors, ) for modeling airspace regulations and physical objectives. As crowd-sourced data is untrustworthy, we assume a uniform translational error on all map features. Table 1 lists the parameters we use throughout our experiments. We draw data from a area, covering Île-de-France, with legal and physical requirements stated in the CoLa program in Listing 4.1. Furthermore, we split the area into tiles where, in each one, two paths from the respectively opposing corners ought to be found, starting and ending at ground level, forming a large-scale network across the metropolitan area.
A Constitution for navigating the skies over Paris: We employ a StaR Map sourced from crowd-sourced OpenStreetMap data to describe a simplified set of airspace regulations applied to the urban environment. Here, while government offices and embassies always require a safety distance for UAVs, the restrictions with regard to the other considered types of map features are altitude, license, and time-dependent. With regard to physical objectives, we consider radio disturbance, noise pollution, risk, and energy consumption alongside legal compliance. {minted} [ frame=none, autogobble, fontsize=, xleftmargin=20pt, linenos ]python # Which StaR Map to use star_map(”./environments/paris.star”).
# Mission parameters describing the circumstances parameter standard_license, expanded_license. parameter daytime, nighttime.
# Limitationss depending on altitude field low_flight_limitations if over(park) or distance(primary) ¡ 30 or distance(secondary) ¡ 15. field mid_flight_limitationss if low_flight_limitations or distance(building) ¡ 20. field high_flight_limitations if mid_flight_limitations or daytime and distance(stadium) ¿ 50 and distance(stadium) ¡ 150.
# Altitude independent government limitations field government_limitations if distance(government) ¿ 200 and distance(embassy) ¿ 200.
# The satisfying airspace over Paris field objective paris_limitations if expanded_license and government_limitations or standard_license and government_limitations and ( altitude ¡ 100 and low_flight_limitations or altitude ¡ 200 and medium_flight_limitations or altitude ¡ 300 and high_flight_limitations ).
# Physical objectives for effective routing field objective radio(”./models/radio.py”). field objective noise(”./models/noise.py”). field objective risk(”./models/risk.py”). path objective energy(”./models/energy.py”).
@@ 4.2. A Constitution for Urban Advanced Aerial Mobility
From OpenStreetMap, we obtain crowd-sourced, real-world data on Paris. To simulate restrictions as they might be imposed on an AAM scenario in urban areas, we consider relations about primary and secondary roads, buildings, sports stadiums, as well as government offices and embassies. Hence, a variety of rules is created that may, in a similar form, be employed to create safe and socially acceptable AAM in densely populated urban areas. Furthermore, in contrast to prior publications (Kohaut et al., 2023a, 2024), we consider altitude-dependent rules for a full three-dimensional airspace regulation.
Figure 2 shows all of the spatial relations and their respective parameters as have been employed for the next steps of UAV routing, probabilistic mission Clearance, Explanation, and Optimization. We choose for the UAV routing setting the parameters daytime and a standard_license, i.e., a comparatively restricted setting in comparison to, e.g., setting the expanded_license parameter instead. The resulting landscapes of probabilities have been visualized in Figure 3.
@@ 4.3. Physical Objectives
First, Figure 4 visualizes the noise objective based on geometry provided by OpenStreetMap. As perceived noise disturbance lessens with an increased altitude of the UAV, so does the cost of the objective function. Second, Figure 5 shows the radio disturbance cost based on data provided by OpenCellID. Here, it is advantageous for the UAV to stay close to at least one of the radio towers to maintain adequate signal quality. The signal quality degrades with the inverse square law as the distance of the UAV to the nearest tower increases. For visualization purposes, the cost of radio disturbance has been normalized across the tiles in the Paris-wide overviews, with the discontinuities stemming from each tile loading the cell tower locations for planning in isolation. Finally, Figure 6 then illustrates the cost function based on the risk of injury. Flying at low altitudes over buildings or water surfaces (compared to roads or parks) reduces this risk. Flying higher increases the risk, as the probability of the crash parabola ending over roads or parks increases. Therefore, the risk map becomes smoother at higher altitudes.
@@ 4.4. Compliant and Effective UAV Routing
Here, we show the resulting networks of paths from VIAS. First, in Figure 7, we show the resulting UAV paths when choosing a sensible trade-off between all of the objectives: noise, radio disturbance, risk, energy, and compliance, respectively. Hence, these paths achieve a balance between physical and social demands. Here, (a) shows a 3D rendering (Google Earth Studio, ) of the network over Paris around the Eiffel Tower, while (b) gives a bird’s-eye overview of the entire covered area with paths drawn in blue if they are likely to satisfy the rules and red if they are likely violating them.
@@
Second, we visualize the extreme-point paths in Figure 8, showing the resulting UAV paths when choosing the extreme points for each of the objectives risk, noise, radio disturbance, energy, and compliance, respectively. Hence, these paths mostly disregard every objective except one in order to best match a single requirement. Furthermore, we show how likely each individual path is by coloring them in blue for highly likely to be compliant paths and red in case they likely violate the rules along the way.
@@
Third, we demonstrate Explanation and Optimization as described in Sections 3.7 and 3.8 in Figure 9. Originally, the parameters in Listing 4.1 were chosen such that the UAV was assumed to be employed during daytime with a standard license and the respectively modeled restrictions, leading to many paths being denied clearance in both the knee-point (Figure 7) and extreme-point paths (Figure 8). By searching the parameter space, we can automatically (i) explain that Clearance is mostly denied due to the choice in license and (ii) propose a more optimal mission setting using the expanded license. Here, we show the rate at which paths obtain clearance with varying thresholds for each extreme point and the knee point paths, once for the original setting (a) and once for the optimal setting (b). In (c), one can see that even the energy-optimal paths are now mostly satisfactory compared to Figure 8.

### v2

@@ 3.2. Statistical Relational Maps for Reasoning on Airspace Regulations
We have recently introduced StaR Maps to encapsulate uncertain environments of semantically annotated features, i.e., geometry with associated stochastic error parameters and labeled with descriptive tags (Flade et al., 2024). Rather than containing a graphical representation of the environment, a StaR Map parameterizes hybrid probabilistic relations between points in state space and sets of features.
@@
According to StaR Maps, a spatial relation with domain and codomain of describes a categorical relationship between a point in , a tag in referencing a set of map features, and a map in .
For example, such a relation may categorize points into whether they lie over any feature with the respective tag. Analogously, a relation with domain and codomain of describes a quantitative relationship, e.g., the distance between a point and the closest element in the set of features with the respective tag. In turn, the set itself depends on the application, e.g., in the UAV routing context, one might have to reference important environment features that an agent may avoid flying over or keep a certain distance to.
@@
As the evaluation of a deterministic spatial relation will differ for each perturbed map, StaR Maps computes the statistical moments to fit the distributions chosen for each relation. To estimate the parameters for a specific spatial relation and a type on a point , we compute . With the set at hand, we obtain the statistical moments, e.g., mean and variance, of .
Using maximum likelihood estimation or moment matching with the desired distribution, for example, a Bernoulli distribution for - or a Gaussian distribution for -domain relations, we store the resulting parameters within a StaR Map. The so-obtained spatial relations can then be compiled into distributional ground atoms in probabilistic first-order logic, which is how we proceed in Section 3.4.
@@ 3.3. Physical Objectives
In practice, agents ought not only to obey traffic laws but also to carefully consider and pursue various physical objectives as well. Here, we consider exemplary physical and geo-referenced objective function formulations evaluating a path regarding the risk of injury for city residents, noise emission, radio signal disturbance, and energy consumption, which have been presented in detail in a prior publication (Hohmann et al., 2023). Hence, in the following, we re-introduce these objective functions briefly and refer to the original work for an in-depth discussion.
We categorize and discuss these objectives in two forms: grid-based and non-grid-based.
@@ 3.4. The Constitutional Language
A Constitution written in CoLa, as exemplified in Listing 2 is written in three parts: First, one defines the setting by referring to a StaR Map, providing the probabilistic spatial relations one may use to encode the legal requirements, and determining the parameters under which the mission may be carried out, e.g., weather conditions, licensing information, or abstract UAV properties. Second, one defines further background knowledge and rules on which the compliance of an agent’s state can be reasoned. Third, one defines which quantities shall be treated as objectives across the navigation space. To this end, CoLa refers to field clauses and objectives if they relate to functions that produce scalar fields across the agent’s state space.
While on the one hand, physical objectives as discussed in Section 3.3 may be incorporated through their mathematical description in a language such as Python 333https://www.python.org/, the logical reasoning on airspace regulations is facilitated by compiling the relevant parts of the Constitution into a probabilistic logic programming language supporting hybrid probabilistic inference, i.e., on discrete and continuous distributions simultaneously, such as Hybrid ProbLog (Nitti et al., 2016). Other systems with more advanced neuro-symbolic capabilities, such as DeepProbLog (Manhaeve et al., 2018), NeurASP (Yang et al., 2020), or SLASH (Skryagin et al., 2023), are left for future work, e.g., to integrate end-to-end learning with deep learning models.
@@
To query for the probability that the Constitution’s modeled requirements are satisfied for a location , it needs to be compiled into the chosen backend, such as Hybrid ProbLog (Nitti et al., 2016). This is achieved by (i) replacing syntactic elements, e.g., conjunctions (and) are written as comma-separated lists in ProbLog, and (ii) inserting the parameters of the StaR Map. We refer to the original paper for details on the latter part (Flade et al., 2024).
Such a probabilistic logic program must first be solved, i.e., the models under which it is satisfied must be found to perform probabilistic inference. While details will differ across off-the-shelf solving pipelines, the rough process can be described in two steps. First, the program is grounded, meaning all variables are substituted, and the program is rewritten in Conjunctive Normal Form (CNF). Second, the actual solver consumes the CNF, producing a Disjunctive Normal Form (DNF) where every conjunction represents a possible assignment of ground atoms under which the (constrained for the currently queried objective) program is satisfied. For instance, one may employ clingo (Gebser et al., 2019), a state-of-the-art tool for solving programmatic first-order logic based on Answer Set Programming (Brewka et al., 2011).
For exact probabilistic inference, one needs to assign the probabilities of atom to take on the value assigned by model . These parameters have been either written down directly in the Constitution, e.g., the take-off mass in Listing 2, or are provided by the StaR Map. In turn, the probability of the Constitution being satisfied given the state and measurement at time is then obtained via the sum-product
@@ 3.5. Many-Objective Path Planning
We facilitate the search for a Pareto set of paths considering the objectives stated in a CoLa program through, e.g., the physical objectives defined in Section 3.3 and airspace-dependent regulations stated in probabilistic first-order logic by employing the Versatile Intelligent Aerial Streets (VIAS) framework for many-objective path planning as presented in a prior publication (Hohmann et al., 2021). Besides the reference implementation of the overall architecture (see Figure 1) being made available as an open-source package, VIAS is also available as a stand-alone solution.
In its first stage, VIAS applies the single-objective Dijkstra algorithm to weighted aggregations of all grid-based objective functions. This results in three-dimensional polygonal paths guaranteed optimal in the employed grid structure. Next, VIAS smooths and approximates the polygonal paths and obtains three-dimensional NURBS curves (Piegl and Tiller, 1995a, b). These splines are henceforth treated as initial solutions in the second stage, a many-objective evolutionary optimization problem calculating an -dimensional Pareto set of smooth paths that can be thought of as a set of optimal trade-off solutions between all considered objective functions.
@@
The smoothing and approximation between the two optimization stages are crucial. At this point, the path representation changes from a polygonal path definition to a NURBS definition. Here, we will only provide a brief discussion; the interested reader is referred to (Hohmann et al., 2023).
VIAS first smoothes the polygonal path by applying a Gaussian filter with kernel . Then, the smoothed paths are approximated in the form of a NURBS curve. For this operation, the critical parameter is the number of control points, , for the resulting NURBS curve. While a low number may lead to an excessive approximation error, a large number increases the size of the optimization vector, expanding the search space during the evolutionary optimization step.
@@ 3.6. Mission Clearance
Given a CoLa program and the subsequently obtained Pareto set of paths from VIAS as outlined in Sections 3.4 and 3.5 respectively, one may still wonder if the results are overall sufficiently likely to satisfy the modeled airspace regulations. Hence, based on our prior publication on Clearance, Explanation, and Optimization in ProMis (Kohaut et al., 2024), we discuss each of these steps with regard to Pareto-optimal proposal paths.
@@
Note that this Clearance does not necessarily coincide with a path being the knee-point or even optimal with regard to the airspace requirements. On the other hand, an extreme point of the Pareto set for an unrelated objective may still be granted Clearance, i.e., this step has to be performed on top of the many-objective routing itself.
@@ 3.7. Mission Explanation
Consider again the example CoLa program in Listing 2. Here, two sets of parameters were defined, one on the employed license and one on the time of day. With regard to granting clearance, it is easy to see in this case that a special operations license allows for easier Clearance since no requirement on visual line-of-sight is imposed.
@@ 3.8. Mission Optimization
| ProMis | Legal Model | Listing 3 | |
@@ 4. Experiments
Throughout Section 3, we have presented an approach for effective and compliant UAV routing under many objectives. In this Section, we evaluate our methods on real-world, crowd-sourced data from the city of Paris, demonstrating how our framework not only jointly models airspace regulations and physical objectives but also produces suitable Pareto sets trading-off the objectives and applying Clearance, Explanation, and Optimization on the mission setting to further improve on the found solutions.
@@ 4.1. Scenario Definition
For our experiments, we consider a large-scale urban environment with data sourced from OpenStreetMap (OpenStreetMap contributors, 2017) and OpenCellID (OpenCellid contributors, 2008) for modeling airspace regulations and physical objectives. As crowd-sourced data is untrustworthy, we assume a uniform translational error on all map features. Table 1 lists the parameters we use throughout our experiments. We draw data from a area, covering Île-de-France, with legal and physical requirements stated in the CoLa program in Listing 3. Furthermore, we split the area into tiles where, in each one, two paths from the respective opposing corners ought to be found, starting and ending at ground level, forming a large-scale network across the metropolitan area.
@@ 4.2. A Constitution for Urban Advanced Aerial Mobility
From OpenStreetMap, we obtain crowd-sourced, real-world data on Paris. To simulate restrictions as they might be imposed on an AAM scenario in urban areas, we consider relations about primary and secondary roads, buildings, sports stadiums, as well as government offices and embassies. Hence, a variety of rules is created that may, in a similar form, be employed to create safe and socially acceptable AAM in densely populated urban areas. Furthermore, in contrast to prior publications (Kohaut et al., 2023a, 2024), we consider altitude-dependent rules for a full three-dimensional airspace regulation. Figure 4 shows all of the spatial relations and their respective parameters as have been employed for the next steps of UAV routing, probabilistic mission Clearance, Explanation, and Optimization. We choose for the UAV routing setting the parameters daytime and a standard_license, i.e., a comparatively restricted setting in comparison to, e.g., setting the expanded_license parameter instead. The resulting landscapes of probabilities have been visualized in Figure 5.
@@ 4.3. Physical Objectives
First, Figure 6 visualizes the noise objective based on the geometry provided by OpenStreetMap. As perceived noise disturbance lessens with an increased altitude of the UAV, so does the cost of the objective function. Second, Figure 7 shows the radio disturbance cost based on data provided by OpenCellID. Here, it is advantageous for the UAV to stay close to at least one of the radio towers to maintain adequate signal quality.
The signal quality degrades with the inverse square law as the distance of the UAV to the nearest tower increases. For visualization purposes, the cost of radio disturbance has been normalized across the tiles in the Paris-wide overviews, with the discontinuities stemming from each tile loading the cell tower locations for planning in isolation. Finally, Figure 8 then illustrates the cost function based on the risk of injury. Flying at low altitudes over buildings or water surfaces (compared to roads or parks) reduces this risk. Flying higher increases the risk, as the probability of the crash parabola ending over roads or parks increases. Therefore, the risk map becomes smoother at higher altitudes.
@@ 4.4. Compliant and Effective UAV Routing
Here, we show the resulting networks of paths from VIAS. First, in Figure 9, we show the resulting UAV paths when choosing a sensible trade-off between all of the objectives: noise, radio disturbance, risk, energy, and compliance, respectively. Hence, these paths achieve a balance between physical and social demands. Here, (a) shows a 3D rendering (Google Earth Studio, 2018) of the network over Paris around the Eiffel Tower, while (b) gives a bird’s-eye overview of the entire covered area with paths drawn in blue if they are likely to satisfy the rules and red if they are likely to violate them.
@@
Second, we visualize the extreme point paths in Figure 10, showing the resulting UAV paths when choosing the extreme points for each objective: risk, noise, radio disturbance, energy, or compliance. Hence, these paths mostly disregard every objective except one to match a single requirement best. Furthermore, we show how likely each individual path is granted clearance by coloring them in blue for a clearance score as defined by Equation 11 and red in case they obtain a low score and likely violate the rules along the way.
@@
Third, we demonstrate Explanation and Optimization as described in Sections 3.7 and 3.8 in Figure 11. Initially, the parameters in Listing 3 were chosen such that the UAV was assumed to be employed during daytime with a standard license and the respectively modeled restrictions, leading to many paths being denied clearance in both the knee-point (Figure 9) and extreme point paths (Figure 10). By searching the parameter space, we can automatically (i) explain that Clearance is mostly denied due to the choice in license and (ii) propose a more optimal mission setting by choosing the expanded license.
In Figure 11, we show the rate at which paths obtain clearance with varying thresholds for each extreme point and the knee point paths, once for the original setting (a) and once for the optimal setting (b). In (c), one can see that even the energy extreme point paths receive mostly high clearance scores compared to Figure 10, meaning that the framework can point the mission design towards settings that allow for satisfactory and highly effective routing even without specifically optimizing paths for compliance.
| Mission Setting | Objective Priority | Area over Curve (Figure 11 (a) and (b)) |
| Standard License | Risk optimal | 0.799 |
| Noise optimal | 0.188 | |
| Radio Disturbance optimal | 0.784 | |
| Energy optimal | 0.791 | |
| Compliance optimal | 0.213 | |
| Balanced (Knee Point) | 0.384 | |
| Expanded License | Risk optimal | 0.108 |
| Noise optimal | 0.056 | |
| Radio Disturbance optimal | 0.112 | |
| Energy optimal | 0.112 | |
| Compliance optimal | 0.037 | |
| Balanced (Knee Point) | 0.075 |
4.5. Runtime Requirements
Our framework aims to provide compliant and effective UAV routing in complex and demanding environments such as urban cores. Because this is a computationally demanding task that depends on the application environment, we discuss our framework’s runtime and current bottlenecks.
In our experiments, we apply our framework to generate two paths for each of the individual airspace volumes across Paris, each spanning of land. Figure 12 shows the time needed for each major step in our pipeline, namely the computation of StaR Map parameters, VIAS’ preparation of physical objectives, and the Dijkstra-based initial search, and VIAS’ evolutionary generation of the Pareto set.
Although the initialization of our framework can be done within minutes, the evolutionary search takes considerably more time and limits the use of many-objective optimization for real-time applications. Instead, we aim to provide a traffic network of compliant and effective UAV paths in an offline fashion. Note how, according to Figure 12, StaR Maps’ runtime is much less dependent on the desired resolution of the search graph.
Instead, as shown in Figure 13, StaR Maps becomes slower with a growing number of map features to consider in the sampling process. Hence, depending on the complexity of the application area, e.g., the number of buildings or roads, StaR Maps computation time may grow considerably.